DIRC Meeting Operations Report
Period: Wednesday 2007/3/28 - Wednesday 2007/4/24
Oncall: Alejandro Perez
Fast Monitoring: Alejandro Perez, Bradley Wogsland, Jose Benitez
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20070424
PEP peak lumi: ~1.0 x 1034
(previous week: ~1.0 x 1034)
BaBar integrated lumi: 2.13 fb-1
(previous week: 3.28 fb-1)
Run 6 PEP delivery: 2.19 fb-1
(previous week: 3.43 fb-1)
Run 6 total delivery: ~30.61 fb-1
Run 6 total BaBar recorded: 29.38 fb-1
including 0 fb-1 off-peak data
PEP activities since this week:
Monday (4/9): VacSealing made on 4/9 near Q4.
Wednesday (4/18): in the morning 2 hours MD for beam characterization.
Thursday (4/19): 12 hours ROD.
Over the weekend (4/20 - 4/23): problems since skew quads installation. Frequent MD studies at low luminosity.
- Brace installed in the support tube, this reduced beam vertical oscillation.
- 12 Decker skew quads istalled (beam focusing).
BaBar activities/issues this week:
4/9: 1 day ROD.
4/19: 12 hours ROD.
- 3/29 Nicolas added a BCS SIAM alarm in the ALH to make sure that we do not forget inhibit when we work on the Chiller.
- Paged 3/31 for high background (scalars all in red) by pilot who claimed MCC was unresponsive to him. Called MCC and they got the scalars back in the green in 15 mins.
- 4/5 New plot for FM (Nicolas) about FX error 11. FX error per ROM
- Since 4/6 the new DFB (replaced on 3/26) put in slot #7, sector #8 does not show up in the calibration plots. This DFB has 16 dead channels, and the mean time for the good channels been shifting. A standalone calibration on 4/9 showed that the timing distributions for all channels of this DFB has lower occupancy (a factor of 10 lower) compared to the other DFBs: still trying to undestand it. In collision data the good channels show ok. DFB7 sector 8
- Since 4/13 power Chiller SIAM glitches. Duration from 0.25sec to 2.0sec occuring at random. Nothing have been done because we had these glitches for a year.
- 4/19: 12 hours ROD,
- Nicolas put an alarm in the ALH about the power DIRC chiller glitches with instructions to the pilot to tell the DQM to go check the DIRC chiller status.
- Chris O'Grady swapped a ROM with a repaired one in order to test it. Got paged at night 4am due to SynchRequest error from the ROM Chris O'Grady swapped. In the morning we had optical fiber disconnection from the same ROM. The ROM was replaced.
- 4/23: Got paged twice in the afternoon because of long periodes of high scaler rates. we ask the pilot to be persistent in complaining MCC. Backgroung Spikes 4/23
logbook for details
- Replace two DFBs, one in sector 0 and one in sector 1 with a group of 8 innefficient or
flaky channels (require bridge).
- Replace new DFB in sector 8 with 16 (TDC chip 0) often dead channels and figure
out what is killing DFBs in that slot. Plot
- Found a X-tree in sector 7 HV group 55, is not very hot and only shows up when taking cosmics.
- DFBs - 8 good, 2 potentially good (one 8 flaky channels, the other configuration problem chich can be bypassed by ODF people), 7 broken
- DCCs - 3 good
- FEE Crates - 2 good
- HV Crates - 1 good shared with DCH
- HV boards - 11 good
Jose, Alejandro and Bradley rotating backup & on-call commissioning